Hunter, Ben, and Blake shared some pencils in the ratio 6:5:1. Blake received 35 fewer pencils than Hunter. How many pencils were there altogether?
Answer by ikleyn(52797)

.
Hunter, Ben, and Blake shared some pencils in the ratio 6:5:1.
Blake received 35 fewer pencils than Hunter. How many pencils were there altogether?
~~~~~~~~~~~~~~~~~

We treat the problem in this way:

    - Hunter has 6x pencils;

    - Ben    has 5x pencils;

    - Blake  has  x pencils.


From the problem,  6x - x = 35,  5x = 35,  which gives  x = 35/5 = 7.


So, the total pencils is  6*7 + 5*7 + 7 = 7*(6+5+1) = 7*12 = 84.    ANSWER

Solved.